OpenStudy (anonymous):

42x^(5)y^(6)z^(10) ___________________ 56x^(3)y^(9) anyone have any clue how to do this?!

OpenStudy (anonymous):

if you're still talking about your restrictions then it's simple you have to make sure that the denominator cannot be zero so x and y cannot be zero those are your restrictions

OpenStudy (anonymous):

i understand that, but i dont know how to solve it

OpenStudy (anonymous):

btw to type fractions you use \frac{numer}{deno} in the equation editor you'll get \frac{2}{4} \[\frac{2}{4} \]

OpenStudy (anonymous):

haha i have no idea how to do that, im not a big computer person i just need to know how to solve this

OpenStudy (anonymous):

well...i dont know how you can solve it... but i can simplify it... you'll get \[\frac{42}{56}*\frac{x^2 * z^{10}}{y^3}\] so actually, the restrictions of the simplified version would be y cannot equal to zero